use definitions or identites to find the exact value of the remaining 5 trig functions.
1. sin=(sqrt3)/4
2. cot=1/2
3. csc=5
4. sec = 5/3
5. cot=2
Thank you so much!!

Answer by Abbey(339)

In order to solve these trig problems, you should remember:
sin^2X+cos^2X=1
and remember that
tan=sin/cos
cot=cos/sin
sec=1/cos
csc=1/sin
1. sin=(sqrt3)/4 then:
cos=(sqrt13)/4
tan=(sqrt3/sqrt13)
csc=4/(sqrt3)
sec=4/(sqrt13)
cot=(sqrt13)/(sqrt3)
2. cot=1/2 then:
sin=2/(sqrt5)
cos=1/(sqrt5)
tan=2
csc=(sqrt5)/2
sec=sqrt5
cot=1/2
3. csc=5 then:
sin=1/5
cos=(2sqrt6)/5
tan=1/(2sqrt6)
sec=5/(2sqrt6)
cot=(2sqrt6)
4. sec=5/3 then:
sin=4/5
cos=3/5
tan=4/3
csc=5/4
cot=3/4
5. cot=2 then:
sin=1/(sqrt5)
cos=2/(sqrt5)
tan=1/2
csc=sqrt5
sec=(sqrt5)/2